You are 35 years old, and have not saved any money yet. You hope to retire at age 65, with a sustainable income of $150,000 per year of current buying power. You assume that inflation will be 3.1% and the fund you want to invest in will return 7.94% per year from now until your death.

a) What is your real rate of return? ______
b) How much money do you need in today's dollars to reach your income goal? $ ______
c) What is that amount in future actual dollars? $ ______
d) how much do you need to save each year to reach your savings goal? $ ______
Answer #1

Answer a:

Real rate of return = Nominal rate of return - Inflation rate = 7.94% - 3.1% = 4.84%

Answer b:

Money do you need in today's dollars to reach your income goal = Required sustainable income per year at current buying power / Real rate of return = $150,000 /4.84% = $3,09,9173

Answer c:

Inflation rate is = 3.1%

Duration up to retirement age = 65 - 35 = 30 years

Future Value Interest Factors for One Dollar Compounded at 3.1% for 30 Periods = 2.498964

Amount in future actual dollars = $$3,099,173 * 2.498964 = $1,270,692

Answer d:

Future Value Interest Factors for a One-Dollar Annuity Compounded at 7.94% Percent for 30 Periods = 127.121

Required saving each year to reach the savings goal = $1,270,692 / 127.121 = $99,95